OpenStudy (anonymous):

What is an equation that is an example of reflecting a function over the x axis?? I will give a medal!!!

OpenStudy (anonymous):

When you reflect a function over the x-axis you negate the function. So, y = x^2. Therefore y = - x^2 is the reflection of y = x^2 over the x-axis.

OpenStudy (anonymous):

You minus the entire function to get a reflection over the x-axis.

OpenStudy (anonymous):

Given: y = 2^x + 2 Reflection over the x-axis: y = -(2^x + 2) or y = -2^x - 2

OpenStudy (anonymous):

ok, thank you! so if I reflect over the y axis it would be like f(x)=(-9x+2)??